OpenStudy (anonymous):

4x-3y=17 2x-5y=-11 then y=?

OpenStudy (callisto):

2x-5y=-11 => 2x= 5y-11 -(1) Put (1) into 4x-3y=17 2( 5y-11) -3y =17 7y = 39 y = 39/7

OpenStudy (anonymous):

39/7 is the final answer?

OpenStudy (callisto):

i think so

OpenStudy (anonymous):

hmm idk man thats not on my worksheet

OpenStudy (callisto):

perhaps you should check the question again to see you you've made some typing mistakes?